### v2

@@ Abstract
As agent systems powered by large language models (LLMs) advance, improving the task performance of an autonomous agent, especially in context understanding, tool usage, and response generation, has become increasingly critical. Although prior studies have advanced the overall design of LLM-based agents, systematic optimization of their internal reasoning and tool-use pipelines remains underexplored. This paper introduces an agent framework grounded in real-world practical experience, with three key innovations: (1) an adaptive prompt generation strategy that aligns with the agent’s state and task goals to improve reliability and robustness; (2) a context-aware tool orchestration module that performs tool categorization, semantic retrieval, and adaptive invocation based on user intent and context; and (3) a layered memory mechanism that integrates session memory, task history, and external summaries to improve relevance and efficiency through dynamic summarization and compression. An end-to-end framework named Jenius-Agent has been integrated with three key optimizations, including tools based on the Model Context Protocol (MCP), file input/output (I/O), and execution feedback. The experiments show a 20 percent improvement in task accuracy, along with a reduced token cost, response latency, and invocation failures. The framework is already deployed in Jenius (https://www.jenius.cn), providing a lightweight and scalable solution for robust, protocol-compatible autonomous agents.
@@ 1 Introduction
Rather than evaluating isolated modules, this paper builds a unified framework named Jenius-Agent, which improves task accuracy, efficiency, and contextual robustness, aligned with emerging agent communication protocols (e.g., MCP, ACP, A2A) IBM BeeAI (2024); Google (2024).
@@
  • We designed a comprehensive evaluation framework covering the procedural, semantic, and efficiency dimensions;
@@ 3.1 Adaptive Prompt Generation
(3) History-aware adaptation: Rather than relying on raw conversation logs, we employs a relevance-driven compression strategy to filter logs, only preserve key cross-turn dependencies. The resulting history is summarized and appended to the prompt, facilitating precise and coherent action planning. For history filtering details, please refer to section of ”Hierarchical Memory Management”.
@@ 3.3 Hierarchical Memory Management
To maintain structural consistency and semantic coherence across turns, this module performs pairwise alignment of adjacent messages in the diaglogue traces. This alignment process detects and rectifies disruptions in the expected message sequence, particularly those arising from partial or failed tool executions (e.g., user-initiated cancellations, API failures, or LLM output errors) that result in missing or empty ToolMessages. In such cases, the system inserts semantically appropriate placeholder entries or reconstructs the absent information through contextual inference. Furthermore, to mitigate data loss caused by system faults or token-length limitations during tool invocation, the mechanism employs a backfilling strategy to recover truncated or omitted messages. By reconstructing these missing components based on available context and execution logs, the module ensures the integrity of the canonical interaction pattern: Human → AI (action proposal) → Tool → AI (response synthesis).
@@ 4.1.1 Execution fidelity
This assessment goes beyond static output quality by examining the rationality and correctness of the agent’s operational trajectory, focus on how it is achieved. As such, it better aligns with the functional requirements of autonomous agents in real-world, dynamic environments where correct behavior depends on both accurate decisions and proper execution order.
@@
Together, TCR, TFR, TIR, and TPS provide a comprehensive assessment of procedural fidelity, capturing both task completion and the rationality of tool usage.
@@ 4.2.1 APIGen: A Single-Turn Tool-Use Benchmark
In our evaluation, we sample 800 instances from APIGen to ensure computational tractability while preserving task diversity. However, a key limitation of APIGen is its restriction to single-turn interactions: each instance represents an isolated tool-calling decision without contextual dependencies or multi-step reasoning. Furthermore, the API descriptions are minimal and syntactic in nature, lacking rich semantic annotations or behavioral specifications. To amplify the challenge and better stress to test our retrieval-augmented architecture, we extend the original candidate tool list for each query by randomly injecting 100 irrelevant tools sampled from its API pool, increasing the average candidate count. This augmented setup intensifies the need for accurate semantic matching and efficient top-$M$ filtering, thereby highlighting the effectiveness of our dynamic tool selection mechanism under high-noise conditions.
@@ 4.4.1 Execution fidelity - 4T testing performances
The results demonstrate distinct patterns between the two datasets. On the public APIGen benchmark, all models perform similarly, with the Base model already reaching a high task completion rate (TCR = 0.8150). Subsequent optimizations (B-P, B-PT, and Jenius) yield only marginal improvements up to 0.8500, reflecting that APIGen mainly assesses pattern matching within familiar tool templates rather than genuine reasoning or generalization.
@@ 4.4.2 Output quality - CRCFF evaluation metrics
We evaluate generated responses using two independent LLM-based evaluators (Qwen and DeepSeek) across five textual quality dimensions: correctness, relevance, completeness, fluency, and faithfulness (CRCFF). As shown in Table 4, the evaluation is conducted solely on Jenius-bench, since APIGen employs virtual API stubs that lack real textual outputs for reliable quality analysis. Jenius-bench, in contrast, features realistic multi-turn dialogues where reasoning traces and tool results jointly determine response quality. The results from both evaluators reveal consistent trends: the Adaptive Prompt Generation module significantly enhances relevance and fluency by producing more stable, context-aware prompts; the Context-Aware Tool Orchestration module improves correctness and completeness through precise semantic alignment and richer tool utilization; and Hierarchical Memory Management strengthens faithfulness by preserving contextual coherence in long-horizon tasks. Overall, the fully integrated Jenius Agent achieves the best CRCFF performance, with an 8–10% improvement in correctness and fluency over the baseline, demonstrating the complementary effectiveness of the three optimization modules in producing coherent, accurate, and faithful responses.
@@ 5.2 Scale and Usage Patterns
  • During the first month of public operation, the system achieved a daily active user (DAU) count of 42 and a monthly active user (MAU) count of 1,277, indicating early but concentrated adoption.
@@ 5.3 Lessons Learned
Real-world deployment of the Jenius Agent provides critical insights into the design and optimization of large language model LLM-based agents.

### v3

@@ Abstract
As agent systems powered by large language models (LLMs) advance, improving performance in context understanding, tool usage, and long-horizon execution has become critical. However, existing agent frameworks and benchmarks provide limited visibility into execution-level behavior, making failures in tool invocation, state tracking, and context management difficult to diagnose. This paper presents Jenius-Agent, a system-level agent framework grounded in real-world deployment experience. It integrates adaptive prompt generation, context-aware tool orchestration, and layered memory mechanism to stabilize execution and improve robustness in long-horizon, tool-augmented tasks. Beyond system design, we introduce an evaluation methodology that jointly measures procedural fidelity, semantic correctness, and efficiency. This framework makes agent behavior observable as a structured execution process and enables systematic analysis of failure modes not captured by output-only metrics. Experiments on Jenius-bench show substantial improvements in task completion rate, with up to a 35% relative gain over the base agent, along with reduced token consumption, response latency, and tool invocation failures. The framework is already deployed in Jenius (https://www.jenius.cn), providing a lightweight and scalable solution for robust, protocol-compatible autonomous agents.
@@ 1 Introduction
Beyond system design, evaluating long-horizon, tool-augmented agents remains underdeveloped. Existing benchmarks emphasize final outputs or single-step tool calls. They fail to capture execution failures such as missing tools, incorrect ordering, or partial completion. This gap motivates an execution-centric evaluation perspective.
@@
Rather than evaluating isolated modules, this paper builds a unified framework named Jenius, which improves task accuracy, efficiency, and contextual robustness, aligned with emerging agent communication protocols (e.g., MCP, ACP, A2A) IBM BeeAI (2024); Google (2024).
@@
  • We introduced a system-level execution abstraction that models agents as structured processes with explicit state, tool interactions, and failure modes.
@@
  • We designed a comprehensive evaluation framework that jointly measures procedural fidelity (4T), semantic quality (CRCFF), and efficiency, enabling diagnosis of real-world agent failures;
@@ 2.3 Memory Management
Prior agent frameworks differ substantially in execution assumptions. For example, AutoGen Wu et al. relies on multi-agent message passing with implicit tool semantics, while ReAct assumes a fixed reasoning–action loop. Memory persistence, tool invocation contracts, and prompt adaptation strategies vary across systems. As a result, direct performance comparison often reflects framework assumptions rather than system effectiveness. Our evaluation therefore focuses on controlled, within-framework evolution to isolate the impact of execution-level design choices.
@@ 3.1 Adaptive Prompt Generation
(3) History-aware adaptation: Rather than relying on raw conversation logs, we employ a relevance-driven compression strategy to filter logs, only preserve key cross-turn dependencies. The resulting history is summarized and appended to the prompt, facilitating precise and coherent action planning. For history filtering details, please refer to section of ”Hierarchical Memory Management”.
@@ 3.2 Context-Aware Tool Orchestration
Tool orchestration in Jenius is not based solely on similarity ranking. All tools are normalized with explicit schemas, including input constraints, expected outputs, and invocation signatures. Retrieval produces a candidate set, but the final tool choice is resolved during execution based on argument compatibility and runtime feedback.
When tools have overlapping scopes or near-duplicate descriptions, multiple candidates may be retrieved. Ambiguity is resolved at execution time: incompatible tools fail to accept arguments or produce invalid outputs, leading to partial execution rather than silent success. These failures propagate through the execution state and affect subsequent planning steps. Similarly, tools with low-quality or adversarial descriptions are more likely to cause invocation errors or incomplete execution. Such cases increase execution inconsistency and reduce procedural correctness, rather than artificially improving apparent success. This behavior is explicitly exposed at the execution level and analyzed in our evaluation, rather than being masked by output-only correctness signals.
@@ 3.3 Hierarchical Memory Management
To maintain structural consistency and semantic coherence across turns, this module performs pairwise alignment of adjacent messages in the dialogue traces. This alignment process detects and rectifies disruptions in the expected message sequence, particularly those arising from partial or failed tool executions (e.g., user-initiated cancellations, API failures, or LLM output errors) that result in missing or empty ToolMessages. In such cases, the system inserts semantically appropriate placeholder entries or reconstructs the absent information through contextual inference. Furthermore, to mitigate data loss caused by system faults or token-length limitations during tool invocation, the mechanism employs a backfilling strategy to recover truncated or omitted messages. By reconstructing these missing components based on available context and execution logs, the module ensures the integrity of the canonical interaction pattern: Human → AI (action proposal) → Tool → AI (response synthesis).
@@ 4.1.1 Execution fidelity
This assessment goes beyond static output quality by examining the rationality and correctness of the agent’s operational trajectory, focusing on how it is achieved. As such, it better aligns with the functional requirements of autonomous agents in real-world, dynamic environments where correct behavior depends on both accurate decisions and proper execution order.
@@
TCR, TFR, and TIR form a complete and mutually exclusive partition of execution outcomes. Their values sum to one for each task. This guarantees full coverage of agent behavior. Improvements in these metrics consistently correlate with higher end-to-end task completion. We therefore use them as diagnostic signals for procedural fidelity rather than universal success criteria.
@@ 4.2.1 APIGen: A Single-Turn Tool-Use Benchmark
In our evaluation, we sample 800 instances from APIGen to ensure computational tractability while preserving task diversity. However, a key limitation of APIGen is its restriction to single-turn interactions: each instance represents an isolated tool-calling decision without contextual dependencies or multi-step reasoning. Furthermore, the API descriptions are minimal and syntactic in nature, lacking rich semantic annotations or behavioral specifications. To amplify the challenge and better stress our retrieval-augmented architecture, we extend the original candidate tool list for each query by randomly injecting 100 irrelevant tools sampled from its API pool, increasing the average candidate count. This augmented setup intensifies the need for accurate semantic matching and efficient top-$M$ filtering, thereby highlighting the effectiveness of our dynamic tool selection mechanism under high-noise conditions. The noise-injected APIGen setting serves as a stress test for retrieval robustness under extreme tool-set expansion. It is not intended to represent realistic tool registries. Our primary conclusions are drawn from Jenius-bench, which reflects real-world tool usage patterns.
@@ 4.2.2 Jenius-bench: A Multi-Turn, Semantically Enriched Task Dialogue Dataset
Jenius-bench is designed for reproducibility. We explicitly exclude time-sensitive or non-deterministic queries, such as date-specific weather or ticket availability. Each task is associated with a fixed tool set and deterministic execution constraints. Reference trajectories specify required tools and ordering, ensuring stable evaluation across runs.
@@ 4.4.1 Execution fidelity - 4T testing performances
The results demonstrate distinct patterns between the two datasets. On the public APIGen benchmark (Table 2), all models perform similarly, with the Base model already reaching a high task completion rate (TCR = 0.8150). Subsequent optimizations (B-P, B-PT, and Jenius) yield only marginal improvements up to 0.8500, reflecting that APIGen mainly assesses pattern matching within familiar tool templates rather than genuine reasoning or generalization.
@@ 4.4.2 Output quality - CRCFF evaluation metrics
We evaluate generated responses using two independent LLM-based evaluators (Qwen and DeepSeek) across CRCFF. As shown in Table 4, the evaluation is conducted solely on Jenius-bench, since APIGen employs virtual API stubs that lack real textual outputs for reliable quality analysis. Jenius-bench, in contrast, features realistic multi-turn dialogues where reasoning traces and tool results jointly determine response quality. In Jenius-bench, session-level summarization is triggered in approximately 32% of tasks. Each task typically spans four to five user–agent turns, with one to two tool invocations per turn. When triggered, summarized segments are compressed to roughly 35–40% of their original token length. This design treats summarization as a regular execution mechanism rather than an exceptional fallback. The results from both evaluators reveal consistent trends: the Adaptive Prompt Generation module significantly enhances relevance and fluency by producing more stable, context-aware prompts; the Context-Aware Tool Orchestration module improves correctness and completeness through precise semantic alignment and richer tool utilization; and Hierarchical Memory Management strengthens faithfulness by preserving contextual coherence in long-horizon tasks. Overall, the fully integrated Jenius achieves the best CRCFF performance, with an 8–10% improvement in correctness and fluency over the baseline, demonstrating the complementary effectiveness of the three optimization modules in producing coherent, accurate, and faithful responses.
@@ 5.2 Scale and Usage Patterns
  • As of now, after six months of experience accumulation and system iteration, the system’s daily active user (DAU) count has increased to 265; in contrast, during the first month of public operation, the system achieved a daily active user (DAU) count of 42 and a monthly active user (MAU) count of 1,277, indicating early but concentrated adoption.
@@ 5.3 Lessons Learned
Real-world deployment of the Jenius Agent provides critical insights into the design and optimization of LLM-based agents.